EDITORIAL

Terry Goodman

We wanted to use the editorial for this isues of MJMS to highlight clarifications of several policies for the journal.

1. The Missouri Journal of Mathematical Sciences publishes well–motivated original research articles as well as expository and survey articles of exceptional quality in mathematical sciences. A section of the MJMS is also devoted to interesting mathematical problems and solutions.

2. Papers should be submitted electronically as a single file to the appropriate associate editors.

3. It is strongly recommended that authors prepare their papers with AMS-TEX, LaTEX, or TEX and files for figures in EPS format.

4. Manuscripts should contain a short abstract, 2000 Mathematics Subject Classification, key words, author’s name, address, email address, and affiliation. Joint authors should indicate on the manuscript which author should receive the proof of the manuscript.

5. Manuscripts should be typewritten in a double–spaced and wide margin format to allow for editorial remarks and should not exceed 10 pages of text.